Mike is 15 years younger than his uncle and 25 years younger than his father. Mike's Great Grandfathers age right now is the sum of the three mentioned previously. If Mike's age right now is 15, hiw old is his Great Grandfather?

Answer by checkley77(12844)

UNCLE IS MIKE +15.
FATHER IS MIKE +25.
MIKE IS 15.
GREAT GRANDFATHER=M+U+F
GG=15+(15+15)+(15+25)
GG=15+30+40=85 YEARS OLD.